A sailor went to three casinos (gambling houses) on the same night. At the Tropicana, he doubled his money and then lost $30; he took his remaining money and went to Ceasar's Palace. There he tripled the money with which he entered and lost $54. At that point, he took his remaining money and went to Aladdins. There he quadrupled the money with which he entered but then spent $72. He spent no more money, and when he returned to his room he had $48 left. How much money did he have before he went to the casinos?: A sailor went to three casinos (gambling houses) on the same night. At the Tropicana, he doubled his money and then lost $30; he took his remaining money and went to Ceasar's Palace. There he tripled the money with which he entered and lost $54. At that point, he took his remaining money and went to Aladdins. There he quadrupled the money with which he entered but then spent $72. He spent no more money, and when he returned to his room he had $48 left. How much money did he have before he went to the casinos?
Answer by jojo14344(1022)


Let x----> amount of money he has before going to the casinos
At Tropicana:
Double the money but lost $30: 2x-30, amount he got
.
At Ceasar's Palace:
Tripled the money and lost $54: 3(2x-30)-54
6x-90-54 ------> 6x-144, amount he got
.
At Aladdins: Quadrupled the money & spent $72: 4(6x-144)-72
24x-576-72 -----> 24x-648, amount he got
.
Back to Room:24x-648=highlight(48), what's remaining.
Continuing by solving x:
24x=648+48=696 --->cross(24)x/cross(24)=cross(696)29/cross(24)
highlight(x=29), Answer ---> Original Money he has before gambling
Let's check by going back,
At Tropicana: 2*29-30=58-30=$28.00
At Ceasar's Palace: 3*28-54=84-54=$30.00
At Aladdins: 4*30-72=120-72=$48.00 Amount he got going back to his room
Thank you,
Jojo
Question 168016: A sailor went to three casinos (gambling houses) on the same night. At the Tropicana, he doubled his money and then lost $30; he took his remaining money and went to Ceasar's Palace. There he tripled the money with which he entered and lost $54. At that point, he took his remaining money and went to Aladdins. There he quadrupled the money with which he entered but then spent $72. He spent no more money, and when he returned to his room he had $48 left. How much money did he have before he went to the casinos?: A sailor went to three casinos (gambling houses) on the same night. At the Tropicana, he doubled his money and then lost $30; he took his remaining money and went to Ceasar's Palace. There he tripled the money with which he entered and lost $54. At that point, he took his remaining money and went to Aladdins. There he quadrupled the money with which he entered but then spent $72. He spent no more money, and when he returned to his room he had $48 left. How much money did he have before he went to the casinos?
Answer by josmiceli(2162)

Let x= the amount he started out with
After the Tropicana, he had
2x - 30 dollars
After Ceasar's Palace, he had
3*(2x - 30) - 54 dollars
After Aladdins, he had
4*(6x - 90 - 54) - 72 dollars
Then he returned to his room with $48
48 = 4*(6x - 90 - 54) - 72
Now solve for x
48 = 24x - 360 - 216 - 72
24x = 360 + 216 + 72 + 48
24x = 696
x = 29
He had $29 before he went to the Casinos
You can go back over it step by step again
He started with $29
After the Tropicana, he had
2*29 - 30 = 28 dollars
After Ceasar's Palace, he had
3*(2*29 - 30) - 54 = 30 dollars
After Aladdins, he had
4*(6*29 - 90 - 54) - 72 = 48 dollars
Then he returned to his room with $48
OK